OpenStudy (anonymous):

MEDAL!!!! What sine function represents an amplitude of 4, a period of pi over 2, no horizontal shift, and a vertical shift of −3? f(x) = −3 sin 4x + 4 f(x) = 4 sin 4x − 3 f(x) = 4 sin pi over 2x − 3 f(x) = −3 sin pi over 2x + 4

OpenStudy (anonymous):

f(x) = 4 sin 4x − 3

OpenStudy (anonymous):

How come ?

OpenStudy (jadeishere):

@mr_basketball27 you have to help them get to the answer >.<

OpenStudy (anonymous):

Thank you @Jadeishere

OpenStudy (anonymous):

The general form is f(x) = Asin(Bx + C) + D (like we did the previous question) Amplitide A = 4 Period = 2*pi / B = pi / 2. Solve for B Phase shift = 0 implies C = 0 Vertical shift D = -3. Plug it into the equation above.

OpenStudy (anonymous):

Ohhh okay thank you Mr. Basketball I am going to tag you in another you could hopefully help me with.